# MMformer with Adaptive Transferable Attention: Advancing Multivariate Time Series Forecasting for Environmental Applications

### v1

@@ Abstract
Environmental crisis remains a global challenge that affects public health and environmental quality. Despite extensive research, accurately forecasting environmental change trends to inform targeted policies and assess prediction efficiency remains elusive. Conventional methods for multivariate time series (MTS) analysis often fail to capture the complex dynamics of environmental change. To address this, we introduce an innovative meta-learning MTS model, MMformer with Adaptive Transferable Multi-head Attention (ATMA), which combines self-attention and meta-learning for enhanced MTS forecasting. Specifically, MMformer is used to model and predict the time series of seven air quality indicators across 331 cities in China from January 2018 to June 2021 and the time series of precipitation and temperature at 2415 monitoring sites during the summer (276 days) from 2012 to 2014, validating the network’s ability to perform and forecast MTS data successfully. Experimental results demonstrate that in these datasets, the MMformer model reaching SOTA outperforms iTransformer, Transformer, and the widely used traditional time series prediction algorithm SARIMAX in the prediction of MTS, reducing by 50% in MSE, 20% in MAE as compared to others in air quality datasets, reducing by 20% in MAPE except SARIMAX. Compared with Transformer and SARIMAX in the climate datasets, MSE, MAE, and MAPE are decreased by 30%, and there is an improvement compared to iTransformer. This approach represents a significant advance in our ability to forecast and respond to dynamic environmental quality challenges in diverse urban and rural environments. Its predictive capabilities provide valuable public health and environmental quality information, informing targeted interventions.
journal: Spatial Statistics
@@
A novel MMformer model incorporating an adaptive transferable multi-head attention mechanism for multivariate time series prediction.
@@
MMformer achieves state-of-the-art (SOTA) performance in multivariate time series, outperforming existing models such as iTransformer, Transformer, and SARIMAX.
@@ 1 Introduction
Deep learning (DL) has become a prevalent approach in multivariate time series (MTS) forecasting, driven by continual advancements in machine learning and its interdisciplinary applications. This progress has significantly enhanced MTS analysis using real-world data, leading to many impactful research contributions [1, 2]. Currently, MTS prediction is at the hot of research, focusing on deep learning-based forecasting, causal analysis, missing value imputation, and integrated heterogeneous MTS modeling [3, 4]. Despite these advancements, MTS analysis faces challenges, particularly related to high-dimensional data noise and complex correlations, which complicate model construction [5]. Additional issues include integrating datasets with varying sampling intervals and missing data [6], enhancing the interpretability of deep learning models [7], and addressing the unique characteristics and objectives of data from diverse domains [8].
MTS analysis primarily utilizes statistical modeling and deep learning methodologies [9], including Generative Adversarial Networks (GANs) [10] and Variational Autoencoders (VAEs) [11], to simulate and model data distributions. Recently, ensemble learning strategies that combine statistical techniques with machine learning algorithms have been proposed, creating hybrid models that use the strengths of both approaches [12]. Traditional statistical models, such as the Generalized Linear Model (GLM) [13], Seasonal Autoregressive Integrated Moving Average (SARIMA) [14], and Autoregressive Integrated Moving Average (ARIMA) [15], are valued for their interpretability and ability to identify linear temporal patterns with minimal data preprocessing. However, their effectiveness is reduced with complex datasets characterized by significant missing values and non-linear patterns. Conversely, machine learning methods like Gradient Boosting Machines (GBM) [16], Random Forests (RF) [17], and Support Vector Machines (SVM) [18] excel at identifying and predicting complex patterns to address challenges of traditional methods to complex data sets with nonlinear patterns but require extensive preprocessing, bringing challenges for users without advanced statistical expertise. Deep learning, a key subset of machine learning, addresses these limitations through its layered architecture and automatic feature learning capabilities. It enables effective non-linear representation and adaptability to diverse data structures to address limitations of traditional methods [19]. However, DL models are often computationally intensive and complex, limiting their practical application [20].
To overcome these challenges, this research introduces MMformer, a novel adaptive framework inspired by the iTransformer [21]. MMformer enhances MTS forecasting by efficiently utilizing temporal characteristics and time dependencies through innovative mathematical methods integrated within the neural network architecture, such as Bayesian inference combined with Meta-Learning. Structure features of MMformer include embedded variate tokens within the Feed Forward Network (FFN) [22] to identify hidden MTS characteristics and Monte Carlo Dropout (MC Dropout) [23] to reduce co-adaptation across network layers, thereby improving model robustness and regularization.
Additionally, MMformer uses an adaptive transferable multi-head attention mechanism, which combines multi-head attention with meta-learning [24]. This optimizes the encoder for various downstream tasks, enhancing the model’s efficiency and generalization capabilities. By effectively capturing temporal information and volatility structures, MMformer improves both predictive performance and processing efficiency.
To validate the effectiveness of MMformer, we apply it to model and predict the Air Quality Index (AQI) [25] and seven key pollutant indices [26] across 331 cities in China from January 2018 to June 2021 and precipitation and temperature patterns at 2415 monitoring sites during the summer (276 days) from 2012 to 2014 [27]. The results demonstrate MMformer’s state-of-the-art ability to handle environmental MTS data, offering enhanced forecasting accuracy and robustness compared to existing models.
@@
    We propose MMformer, an innovative approach that integrates the dimension transpose with innovative Adaptive Transferable Multi-Head Attention (ATMA) and Bayesian approximate inference to better identify latent features in MTS data. This integration minimizes information loss and enhances representation learning, improving the ability to capture temporal patterns and forecasting accuracy.
@@
    We develop an advanced adaptive transferable multi-head attention module incorporating Model-Agnostic Meta-Learning (MAML). This design optimizes the attention mechanism, boosts model performance, reduces the need for manual tuning, and saves time.
@@
    Our MMformer is the first to demonstrate superior capability in long-term forecasting across various locations and multiple variables, particularly in predicting environmental issues, which is state-of-the-art performance. It also assists actuaries in optimizing insurance strategies for respiratory conditions and catastrophe insurance and supports government ecological policy-making.
@@
    Empirical analysis demonstrates that in air quality datasets, our approach significantly enhances predictive performance, reducing Mean Squared Error (MSE) by 50%, Mean Absolute Error (MAE) by 20%, and relative to other models except SARIMAX a 20% decrease in Mean Absolute Percentage Error (MAPE), our model outperforms Transformer and SARIMAX, achieving approximately 30% reduction in MSE, MAE, and MAPE metrics. Furthermore, better performance was observed in comparison with iTransformer. These findings substantiate the hypothesis that model efficacy correlates positively with feature dimensionality, suggesting that higher-dimensional feature spaces enable more robust predictive capabilities in MMformer.
The structure of this paper is organized as follows: Section 2 reviews MTS forecasting methods. Section 3 details the proposed approach. Section 4 presents the experimental design and results. Finally, Section 5 concludes and discusses future work.
@@ 2 Related Work
Advances in neural network architectures, particularly Transformer and its variants, have significantly elevated the prominence of multivariate time series (MTS) forecasting methodologies, catalyzing multidisciplinary research collaborations and scholarly investigations across diverse domains [28, 29, 30, 31, 32]. This section presents a systematic review of the evolution of MTS forecasting techniques, tracing their development from conventional statistical approaches to contemporary neural network-based methods, thereby establishing the theoretical foundation for our proposed MMformer model.
2.1 Traditional Time Series Prediction
Early efforts in MTS forecasting mainly relied on statistical models such as Autoregressive Integrated Moving Average (ARIMA) [33] and Seasonal ARIMA (SARIMA) [34]. These models are experts in capturing linear dependencies in univariate time series data but face significant limitations when extended to multivariate settings [35, 36, 37]. Specifically, ARIMA and SARIMA struggle with the increased complexity and computational demands of modeling multiple interrelated variables [38]. Furthermore, these models operate under the restrictive stationary assumption, requiring the statistical properties of the time series to remain constant over time, which is often violated in real-world multivariate data [box_time_2015]. Similarly, Exponential Smoothing [33] is simple and offers interpretability but is constrained by high-dimensional data, relies on stationary or locally stationary processes, and fails to capture nonlinear relationships effectively [39].
2.2 Multivariate Time Series Prediction
To address the limitations of univariate models, multivariate approaches like Vector Autoregression (VAR) [40] and State Space Models (SSMs) [41] have been introduced. VAR models extend the autoregressive framework to multiple variables, allowing for the capture of interdependencies among different MTS [42]. Despite their theoretical appeal, VAR encounters challenges related to high dimensionality, significantly as the number of variables increases, leading to computational inefficiency and overfitting [43]. While flexible in representing dynamic systems, SSMs often require extensive computational resources and expertise to specify appropriate latent structures, making them less accessible for large-scale applications [44].
2.3 Machine Learning in Time Series Prediction
The advent of machine learning techniques brought new perspectives to MTS forecasting [12]. Support Vector Machines (SVMs) [45] and the Multilayer Perceptron (MLP) [46] as popular choices due to their ability to model complex, nonlinear relationships without extensive parameter tuning. These methods improved prediction accuracy compared to traditional statistical models [47]. However, they lack mechanisms to capture temporal dependencies effectively, necessary for MTS data [47, 48]. Additionally, feature engineering becomes a prerequisite to harness temporal information, adding to the complexity of model development [12].
2.4 Deep Learning in Time Series Prediction
Recent advancements in MTS analysis, with more researchers moving away from traditional MTS forecasting to explore machine learning and neural network methodologies [35, 49, 50].
Deep learning models, particularly Recurrent Neural Networks (RNNs) [51] and Long Short-Term Memory networks (LSTMs) [52], changed MTS forecasting by modeling temporal dependencies. RNNs introduced the capability to retain information over sequences [53], while LSTMs reduced the vanishing gradient problem, enabling the capture of long-range dependencies [54]. Despite these advancements, RNN-based models require many computational resources and training time [6]. Furthermore, they often struggle with parallelization, limiting their scalability for large datasets and real-time forecasting applications [55, 56].
2.5 Transformer in Time Series Prediction
The introduction of Transformer architectures represented a substantial advancement in MTS forecasting [57]. Models such as the Temporal Fusion Transformer (TFT) [58] and Informer [59] leveraged self-attention mechanisms to capture dependencies across different time steps efficiently [28]. These models offer advantages in handling long-range dependencies and parallel processing, resulting in faster training times and improved scalability [57]. However, challenges remain in optimizing self-attention for MTS and effectively integrating diverse data modalities to enhance prediction accuracy [60, 61].
2.6 Structured Diversity in Time Series Prediction
While simple and interpretable, traditional MTS models often struggle with nonlinear and complex data [62]. In contrast, neural networks offer improved predictions for such data but require extensive training and careful parameter selection to avoid overfitting [35]. Integrating structural diversity allows models to capture a broader range of patterns and interactions within the data, reduce individual model biases, and improve robustness [63]. However, for complex, real-world datasets, neural networks are often the preferred forecasting tool [64, 65]. Despite these benefits, effectively combining diverse structures remains complex, requiring complicated architectures and extensive computational resources.
2.7 MMformer: Advancing Multivariate Time Series Prediction
Based on the concept in the iTransformer [21] and addressing the Transformer’s limitations in time series prediction, the proposed MMformer introduces a novel approach to MTS forecasting. Meta-learning helps enhance time series forecasting when data is limited [66]. The Meta-Learning-based prediction mechanism for few-shot time series forecasting demonstrates superior accuracy and convergence speed, surpassing models trained directly on targeted task-specific MTS datasets [67]. Therefore, MMformer integrates the adaptive transferable multi-head attention with the iTransformer’s dimension transpose for high-dimensional, nonlinear data, incorporating spatial dependencies alongside temporal dynamics. By effectively capturing multiple data patterns and using structural diversity, MMformer enhances prediction accuracy and scalability. Furthermore, it optimizes computational efficiency, making it suitable for real-time applications. Through these innovations, MMformer addresses the challenges posed by traditional and contemporary models, offering a robust solution for complex air quality forecasting tasks.

### v2

# MMformer with Adaptive Attention: Advancing Multivariate Time Series Forecasting for Environmental Applications

@@ Abstract
Multivariate time series forecasting is critical in numerous scientific and engineering fields. Current deep learning models, however, face fundamental challenges in this domain. They struggle to handle dynamically changing data distributions, achieve model self-adaptation, and quantify prediction uncertainty. These limitations are particularly severe in fields like environmental science, where high-reliability decisions are essential. To address these limitations, we introduce MMformer, an innovative common adaptive multivariate time series forecasting model. MMformer’s core lies in the synergistic integration of three key components: an encoder-only optimized architecture for capturing temporal dependencies; the Adaptive Transferable Multi-Head Attention mechanism, using meta-learning to enhance model adaptability and generalization; and Monte Carlo Dropout for improved robustness and crucial uncertainty quantification.
We rigorously evaluate MMformer on real-world environmental datasets (e.g., 1277 days of air quality data across 331 Chinese cities, 1826 days of temperature and rainfall data from 909 stations) and general MTS benchmarks (PEMS). Results demonstrate that MMformer achieves superior performance in accuracy, adaptability, and uncertainty quantification than iTransformer, PatchTST, TimesNet, and Transformer. Specifically, on air quality data, MMformer reduced MSE, MAE, and MAPE by up to 70.195%, 37.240%, and 36.542%, respectively, compared to iTransformer; up to 69.641%, 37.121%, and 35.632% compared to TimesNet; up to 71.578%, 40.288%, and 39.383% compared to PatchTST; and up to 68.182%, 36.641%, and 35.343% compared to Transformer. On climate data, the performance improvements mirror those on the air quality dataset. MMformer also achieved optimal results on PEMS03, PEMS04, and PEMS08, demonstrating strong generalization. Our approach offers key theoretical insights for researchers and significant practical value for policymakers. It marks a significant advance in forecasting and managing dynamic environmental challenges. This capability enables precise interventions to protect public health and promote environmental sustainability.
journal: Applied Soft Computing
@@
A novel MMformer model with a novel adaptive forecasting framework for multivariate time series prediction.
@@
MMformer achieves state-of-the-art performance on widely-used MTS benchmarks and challenging real-world environmental prediction tasks, outperforming existing models such as iTransformer, PatchTST, TimesNet, and Transformer.
@@ 1 Introduction
Deep learning (DL) leads multivariate time series (MTS) forecasting, propelled by machine learning advancements and broad interdisciplinary adoption. Current advanced models, however, face significant limitations. They struggle with dynamic data distributions, model self-adaptation, and quantifying prediction uncertainty [1, 2, 3, 4]. These challenges are particularly critical in high-risk decision-making domains like environmental science [5, 6]. Forecast reliability directly impacts public health and sustainable development strategies [7]. Reliable uncertainty estimation is crucial for effective environmental decision-making [8]. Standard deep learning architectures frequently lack this vital capability. This study employs advanced soft computing techniques to address these pervasive MTS prediction challenges. We validate their effectiveness in demanding environmental applications.
MTS analysis has evolved from traditional statistical models to DL [9]. Traditional models, like Generalized Linear Models (GLMs) [10], Seasonal Autoregressive Integrated Moving Average (SARIMA) [11], and Autoregressive Integrated Moving Average (ARIMA) [12], offer interpretability but struggle with the inherent complexity, nonlinearity, and high dimensionality of modern MTS data. Conversely, DL methods excel at automatic feature extraction and modeling complex data [13]. The above motivates the development of our innovative DL model with three motivations: (1) enhancing the utility of the predictions, particularly in environmental, through developing robust models and uncertainty analysis; (2) improving prediction reliability and informing environmental decision-making through uncertainty quantification; and (3) improving predictive accuracy, optimizing resource management, and facilitating the understanding of environmental changes by effectively handling spatiotemporal data. However, three key challenges persist: (1) standard DL models can be computationally intensive, less interpretable, and lack uncertainty analysis [14]; (2) achieving the adaptability and robustness for reliable uncertainty quantification across diverse real-world conditions remains challenging [15]; and (3) the predictive performance of existing DL models on spatiotemporal time series is often suboptimal [16]. A performance bottleneck persists for common DL models, which have limited accuracy with complex spatiotemporal dependencies and inter-variable interactions in real-world data [17]. These persistent challenges necessitate an innovative DL model to enhance MTS prediction adaptability, robustness, and reliability systematically.
To overcome these challenges, this study introduces MMformer, a novel adaptive integrated forecasting model. The core innovation of MMformer lies in the synergistic integration of three key components within a neural network architecture, specifically designed to enhance adaptability and provide uncertainty quantification. These components are: (1) an Encoder-only inverse embedding neural network for capturing temporal dependencies [18]; (2) an Adaptive Transferable Multi-head Attention (ATMA) mechanism that incorporates meta-learning strategies [19], specifically chosen to significantly enhance the model’s adaptability and generalization across multiple regions and data distributions, thereby addressing the need for domain-specific solutions; and (3) Monte Carlo Dropout (MC Dropout) [20], strategically applied to improve regularization, model robustness against noise, and crucially, to provide vital uncertainty quantification for better risk assessment. The integration of meta-learning addresses the need for effective adaptation, and MC Dropout enhances interpretability through uncertainty modeling, directly responding to identified motivations for model design.
We evaluate MMformer’s performance against state-of-the-art (SOTA) and baseline models, including iTransformer [18], PatchTST [21], TimesNet [22], and Transformer [23], across diverse datasets. The evaluation encompasses the PEMS datasets [24], and two challenged real-world environmental datasets: (1) seven major pollutants across 331 Chinese cities (January 2018 - June 2021, 1277 days) [25], and (2) precipitation and temperature data from 909 sites (2018-2022, 1826 days) [26]. Validation on these real-world datasets, alongside the widely used MTS dataset, addresses concerns about domain specificity and dataset validity. The comprehensive comparisons demonstrate significant performance gains, proving the model’s effectiveness.
@@
    We introduce MMformer, an innovative model designed for the specific challenges of MTS forecasting across multiple regions. By integrating an encoder-only network, the ATMA module, and MC Dropout, MMformer effectively captures complex time features, generalizes across diverse conditions, and provides robust uncertainty quantification. The design directly addresses the limitations of DL models in risk assessment and decision-making.
@@
    We developed the ATMA module, which optimizes the attention mechanism and improves overall model performance. ATMA enables rapid adaptation to different regions’ prediction tasks, enhances data heterogeneity and dynamics robustness, and reduces reliance on domain-specific knowledge. By performing gradient updates on each sub-task, ATMA facilitates fast adaptation and improves generalization across tasks via meta-gradients, thereby contributing to overall model performance and generalizing across different data distributions and varying regions.
@@
    We incorporate MC Dropout to enhance model robustness in noisy environments and provide prediction uncertainty quantification, enhancing result interpretability and supporting risk assessment and decision-making.
@@
    Empirical results demonstrate MMformer’s superior performance in MTS prediction, particularly in forecasting for multiple regions. On air quality datasets, MMformer achieves significant improvements, reducing MSE by 68.182%–71.578%, MAE by 36.641%–40.288%, and MAPE by 35.343%–39.383% compared to iTransformer, Transformer, TimesNet, and PatchTST. MMformer also exhibited excellent performance on climate datasets, with all evaluation metrics decreasing by 27.837%–55.363% relative to baselines. Experiments on the PEMS datasets further validate MMformer’s generalization capability. On PEMS03, MMformer achieves the best performance, with all evaluation metrics decreasing by 3.352%–43.103% compared to baselines. On PEMS04, MMformer also demonstrated the best performance, reducing all evaluation metrics by 10.714%-39.458%. On PEMS07, MMformer’s MSE is comparable to TimesNet but underperformed iTransformer; however, it still improves all metrics compared to PatchTST and Transformer, reducing by 7.143%-54.819%. On PEMS08, MMformer shows the best performance, reducing all evaluation metrics by 10.672% - 69.470%. The above experimental outcomes underscore MMformer’s strong generalization ability and robustness in complex multiple regions MTS forecasting.
The structure of this paper is organized as follows: Section 2 reviews MTS forecasting methods. Section 3 details the proposed approach. Section 4 presents the experimental design and results. Section 5 discusses the research objectives and the research results in relation to previous work. Finally, Section 6 concludes and discusses future work.
@@ 2 Related Work
MTS forecasting is critical in fields such as environmental science. Methodologies have advanced significantly with the rapid development of neural network architectures, notably Transformers and their variants [27, 28, 29, 30, 31]. While traditional statistical methods (e.g., ARIMA, SARIMA) and early machine learning techniques (e.g., VAR [32], SVMs [33], Multi-Layer Perceptrons (MLPs) [34]) once dominated time series analysis, they exhibit limitations when addressing the inherent complexities of real-world MTS data. These limitations include handling high dimensionality, intricate nonlinear relationships, spatio-temporal dynamics, cross-domain adaptability, and capturing subtle inter-series dependencies and long-range dependencies. Consequently, research focus has shifted towards DL models to more effectively tackle these key MTS forecasting challenges. This section will detail the evolution of DL in MTS analysis, emphasizing how these models overcome challenges and discussing current progress in model robustness and generalization, thereby laying the theoretical groundwork for the proposed MMformer.
2.1 Early Deep Learning Models for MTS Forecasting: RNNs and LSTMs
Deep learning has driven the evolution of MTS forecasting, transitioning from traditional statistical and early machine learning methods to neural network-based approaches. The shift is underpinned by deep learning’s superior capacity to capture intrinsic temporal dependencies [35, 36] and its reduction in the need for extensive manual feature engineering [37, 38, 39]. Recurrent Neural Networks (RNNs) [40], including variants like Long Short-Term Memory (LSTM) [41] and Gated Recurrent Unit (GRU) [42], significantly advanced MTS prediction by modeling temporal dependencies. RNNs introduced the capability for information retention across time steps [43], while LSTMs and GRUs mitigated the vanishing gradient problem, enabling the capture of long-range dependencies [44, 43]. However, RNN-based models exhibit inherent limitations. Their sequential processing architecture hinders parallel computation efficiency for long sequences, leading to increased training times and resource consumption [2]. They also face challenges in capturing complex inter-variable interactions [2] and, due to their serial nature, are difficult to parallelize, thereby limiting scalability for large datasets and real-time applications [45, 46].
2.2 Beyond RNNs: Attention Mechanisms and Transformer Models in MTS Prediction
Attention mechanisms represent a significant advancement in MTS forecasting, overcoming RNN limitations by enabling models to selectively weight input sequence elements and thus better capture long-range dependencies [47]. The Transformer architecture, built upon self-attention, directly computes dependencies between any two sequence positions. Attention eliminates reliance on sequential order, greatly enhancing the ability to capture long-range temporal dependencies [23]. Furthermore, self-attention’s inherent parallelizability is crucial for processing high-dimensional, long-span MTS data.
Transformer architectures and their variants have been extensively developed for MTS forecasting, enhancing their capability to model complex dynamics [27]. Informer reduced computational complexity via Probsparse self-attention and Auto-correlation mechanisms, facilitating efficient processing of ultra-long sequences [48]. PatchTST segments time series into patches, applying Transformers to capture both local intra-patch and global inter-patch dependencies, achieving strong benchmark performance [21]. TimesNet focuses on time series decomposition, separating trend and seasonality to better model temporal patterns across various scales, resulting in SOTA performance [22]. Notably, iTransformer reorients attention to the variable dimension, directly addressing the critical task of capturing inter-variable correlations in MTS [18].
These attention-based and Transformer models excel at capturing complex nonlinearities, spatio-temporal dynamics, and subtle inter-variable dependencies in MTS data through their robust feature extraction and deep sequence modeling [23]. iTransformer’s specific focus on inter-variable correlations directly addresses a core MTS challenge. Nevertheless, standard Transformers retain a quadratic computational complexity issue when processing extremely long sequences. Ongoing research thus focuses on novel optimization strategies to improve precision, robustness, and adaptability to diverse data scales and heterogeneities [49, 50].
2.3 Uncertainty Quantification, Adaptability, and Meta-Learning in Deep Learning for MTS Forecasting
While Transformers have significantly advanced MTS forecasting, particularly in handling long-range dependencies and enabling parallel processing, the complexity of real-world MTS data, especially in applications across multiple regions, necessitates further exploration [51, 52]. It includes capturing finer aspects of data dynamics, such as inherent predictive uncertainty and model adaptability, crucial for reliable forecasts and informed decision-making [53].
Probabilistic forecasting and uncertainty quantification are vital for high-stakes applications like environmental monitoring, where understanding forecast uncertainty is as important as the prediction itself [54]. Advanced deep learning methods provide mechanisms to quantify uncertainty beyond point predictions. Techniques like MC Dropout and Bayesian neural networks enable models to express confidence levels in their forecasts [20, 55]. MC Dropout, in particular, offers a practical approach to estimate prediction variance, a capability that deterministic models typically lack. Uncertainty quantification is essential for risk assessment and decision-making in dynamic environments.
Real-world MTS datasets often exhibit significant statistical property variations across different tasks or time periods [56]. Meta-learning and transfer learning have emerged as key solutions to enhance model adaptability and generalization [57]. These methods enable models to "learn to learn" by acquiring knowledge from task distributions. It facilitates rapid adaptation to novel or unseen MTS forecasting challenges, even with limited target task data [58]. The focus on rapid adaptation directly addresses the heterogeneity prevalent in environmental data.
Through these advanced deep learning techniques, researchers are developing more robust, adaptable, and interpretable MTS forecasting models to address the complex challenges posed by real-world spatiotemporal time series data.
2.4 MMformer: A Meta-Learning Framework for Adaptive MTS Forecasting
MTS prediction, particularly in environmental science, faces critical challenges in model adaptability, cross-task generalization, and quantifying prediction uncertainty. While Transformer architectures advance long-range dependency capture and parallel computation, standard models are limited by high-dimensional MTS data, complex inter-variable relationships, and data heterogeneity. To address these, MMformer is proposed, building on prior work, notably iTransformer’s dimension inversion approach, to resolve key MTS prediction challenges.
MMformer’s core innovation is its ATMA mechanism, integrating Meta-Learning for rapid parameter adaptation and optimization when facing data distribution shifts or new tasks. The capability addresses the critical need for domain adaptability across multiple regions of MTS data with varying conditions. Furthermore, MMformer’s embedding strategy, which has timestamps and transposes dimensions, enhances its capture of intrinsic spatio-temporal dynamics.
To bolster prediction reliability, MMformer strategically incorporates MC Dropout. MC Dropout enhances robustness against noise and quantifies uncertainty, vital for risk assessment and decision-making where deterministic models fall short. By effectively capturing diverse data patterns and leveraging its structural diversity, MMformer improves prediction accuracy, scalability, and robustness for real-time applications. These integrated innovations offer a more robust and intelligent solution for complex multi-regional MTS forecasting, overcoming previous limitations.